# A Resolution Independent Neural Operator

### v1

@@ 2.2 Dictionary Learning
We have studied the ability of the proposed algorithm to find appropriate orthogonal basis functions in Appendix E, where we compare our algorithm with classical Gappy PCA or Gappy Proper Orthogonal Decomposition (GPOD) [11] in identifying basis functions from masked, incomplete data.
@@ 3 Numerical Examples
In this section, we showcase the resolution independence of RINO by solving four numerical examples, where the data is synthesized from PDE solutions under different setups, ranging from 1D operators to operators with different supports in their input and output functions. The setup for the data generation and operator learning task are specified for each problem seperately in Sections 3.1, 3.2, 3.3 and 3.4.
In cases where the PDE solver for data generation works on a fixed discretization, we randomly select only a subset of input function sensors for training to showcase RINO’s ability to handle irregular data. The locations and numbers of these sensors vary from sample to sample. Assuming an input function realization is discretized at $M$ points, the “random” data is organized by randomly selecting $M_{\text{rand}}$ points from these, where $M_{\text{rand}}$ is drawn from the uniform distribution $M_{\text{rand}}\sim\text{Uniform}[M_{\text{min}},M_{\text{max}}]$, and $0<M_{\text{min}}<M_{\text{max}}\leq M$.
@@
3.1 Example 1: Antiderivative
@@
where $u(x)$ is modeled as a Gaussian random process with zero mean and a covariance function defined by the radial basis function (RBF) with length scale $l=0.2$, i.e., $u(x)\sim\mathcal{GP}(0,\text{Cov}(x_{1},x_{2}))$; details are provided in Appendix F. We use the open-source dataset from [17], which includes 150 pairs of input-output functions for the training set and 1000 pairs for the test set. In the original dataset, the functions are discretized at 100 equally spaced sensor points. Following the procedure described earlier, we set $M_{\text{min}}=10$ and $M_{\text{max}}=60$ to generate random point clouds for the training and testing cases.
@@
As described in the formulation section, the first step in RINO is to learn a dictionary for representing input functions. This dictionary will be used to find embedding coordinates for the operator learning task, specifically for use in the branch network. Figure 3(a) shows the reconstruction errors in Algorithm 1 during the ADAM [15] iterations. Each red cross on the curve indicates the step at which a new basis function is added (line 8 in Algorithm 1). With a total of 9 learned basis functions, the data can be reconstructed fairly accurately, as shown by the distribution of errors in Figure 3(b) for both training and testing datasets. The learned basis functions are plotted in Figure 3(c).
@@
Figure 4: Antiderivative Example: (a) Loss function during optimization iterations for learning the operator. (b) Distribution of output function prediction errors for the training and testing datasets after training.
@@
The ADAM iterations for operator learning using the input function embeddings obtained from the learned dictionary are shown in Figure 4(a). The close gap between the training and testing errors during the iterations indicates good generalization behavior, suggesting that the input function representation is well informative. In Figure 4(b), the distribution of output function prediction errors after training is shown.
@@
Figure 5: Antiderivative Example: Prediction of the RINO for two queried input functions at inference and for the testing dataset. (a) Shows the sample with the highest error in output function prediction, and (b) shows the sample at the 25th percentile of the error distribution (i.e., 75% of cases have less error than this value). The first column shows the queried input function data, the second column shows the projection coefficients of the queried input function data onto the learned dictionary, see Figure 3(c), and the last column shows the output function prediction.
@@
In Figure 5, the prediction of RINO at inference is shown for two queries from the test dataset. Figure 5(a) presents the results for the worst-case output function prediction. The first column shows the queried input function, which includes a point cloud with 35 data points. The second column shows the projection coefficients of the input function data onto the learned dictionary, which has a total of 9 basis functions. The reconstruction using these coefficients is also plotted in the first column figure. The last column shows the prediction of the output function. Similarly, Figure 5(b) presents the same results for the test sample at the 25th percentile of the error distribution, see Figure 4(b), meaning 75% of test cases have less error than this sample. As the results show, the RINO method demonstrates reasonably good performance in providing predictions at inference for input function data recorded at an arbitrary (but sufficiently rich) number of sensors and locations.
3.2 Example 2: Nonlinear 1D Darcy’s Equation
@@
The worst-case predictions of RINO are shown in Figure 6(a), demonstrating visually acceptable accuracy for both the output function prediction and the input function reconstruction. Similarly the test sample at the 25th percentile of the error distribution is shown in Figure 6(b). The sparse coding feature of the dictionary learning is evident from the second columns, where the contributions of many basis functions are negligible.
Figure 6: 1D Nonlinear Darcy’s Equation: Prediction of the RINO for two queried input functions at inference and for the testing dataset. (a) Shows the sample with the highest error in output function prediction, and (b) shows the sample at the 25th percentile of the error distribution (i.e., 75% of cases have less error than this value). The first column shows the queried input function data, the second column shows the projection coefficients of the queried input function data onto the learned dictionary, and the last column shows the output function prediction.
3.3 Example 3: Nonlinear 2D Darcy’s Equation
@@
where the solution-dependent permeability is $\kappa(s(\boldsymbol{x}))=0.2+s^{2}(\boldsymbol{x})$ and the source term is a random field $u(\boldsymbol{x})\sim\mathcal{GP}$ with length scale $l=0.25$. Homogeneous Dirichlet boundary conditions $s=0$ are defined at the domain boundaries. The solution field is obtained via the finite element method. Input function random fields are discretized on a $20\times 20$ uniform grid, while the triangulation of the finite element domain does not necessarily coincide with these input sensor point locations. Following the procedure described earlier, we set $M_{\text{min}}=100$ and $M_{\text{max}}=280$ to generate random point clouds for the training and testing cases. The sizes of the training and test datasets are 800 and 200, respectively.
The predictions corresponding to the worst-case test error for the output field is shown in Figure 7. While RINO’s accuracy is relatively satisfactory, it could be improved by conducting a rigorous hyperparameter search to find a better neural network configuration.
Figure 7: 2D Nonlinear Darcy’s Equation: Worst-case output function prediction $s(\boldsymbol{y})$ of RINO in the test dataset. The top left figure shows the queried input function data $u(\boldsymbol{x})$, where the red cross symbols indicate the masked information. The middle top figure shows the input field reconstruction via the learned dictionary. The top right figure shows the error between the reconstructed input field and the ground truth. The bottom figures show the output field $s(\boldsymbol{y})$ prediction and its corresponding errors.
3.4 Example 4: Burgers’ Equation
@@
where periodic boundary conditions are assumed, and the periodic initial condition is a Gaussian random process with parameters similar to [35], and $\nu=0.01$. A spectral solver is used to obtain the solution field on a $101\times 100$ grid in the space-time domain for each input function, which is discretized on 101 uniformly spaced points. Following the procedure described earlier, we set $M_{\text{min}}=40$ and $M_{\text{max}}=70$ to generate random point clouds for the training and testing cases. The sizes of the training and test datasets are 1500 and 500, respectively. Notice that the input function samples are scattered and randomly picked; however, the predictions are on the full solution grid for the results in this section, similar to other examples.
@@
To showcase the flexibility of the RINO framework, we use predefined trunk basis vectors as the first 70 POD bases obtained from the solution output snapshots of the training dataset. It has been shown previously that using predefined optimal bases can facilitate the training of DeepONet and achieve faster convergence [18]. Two samples from the test cases are shown in Figure 8.
@@
Figure 8: 2D Burger’s Equation: Prediction of the RINO for two queried initial condition fields at inference and for the testing dataset. (a) Shows the sample with the highest error in output function prediction $s(\boldsymbol{y})$, and (b) shows the sample at the 25th percentile of the error distribution (i.e., 75% of cases have less error than this value). The first column shows the ground truth solution, the second column shows prediction, and the last column shows the point-wise error. The output domain $\boldsymbol{y}=(x,t)$ has spatio-temporal coordinates, where the horizontal axis represents space and the vertical axis represents time. The red cross symbols in the middle column indicate the masked locations of the queried initial condition.
3.5 Discussion
@@
After training the models on those randomly sampled data, we can perform inference during testing using any combination of input function sensors and at any desired points, thanks to the fully continuous nature of our method for processing both input and output functions. To examine the influence of discretization, we also perform some test cases on regularly subsampled sensors for the input function at inference. A summary of our results is provided in Table 1. Rows in this table with a numeric value for the number of sensors (e.g., $M=100$) indicate performance during inference for input functions queried at structured and uniformly spaced sensor locations.
@@
The error statistics shown in Table 1 are computed for five trained models. Each neural network model is randomly initialized before training, and the sensor locations for the input functions are also randomly generated for each training run.
@@
As the results suggest, the output function prediction is nearly insensitive to the discretization resolution of the queried input function for all PDE cases. However, in the 1D Nonlinear Darcy’s and Burgers’ equations, we observe a degradation in accuracy after some level of resolution coarsening (see the last row for each case). This occurs because, as anticipated, when the resolution falls below a certain threshold—particularly when it is significantly less than the problem’s characteristic length scale—the predictions become erroneous and thus resolution-dependent.
Number of Sensors Test Error (mean $\pm$ std) Train Error
@@
Table 1: Resolution dependence study for several datasets by reporting the output function prediction errors for queried input functions at various resolutions.

### v2

@@ 2.2.1 Batch-wise Learning
We have studied the ability of the proposed algorithm to find appropriate orthogonal basis functions in Appendix F, where we compare our algorithm with classical Gappy PCA or GPOD [13] in identifying basis functions from masked, incomplete data. Moreover, the possibility of using random basis functions is also explored and discussed.
2.2.2 Sample-wise Learning
By assuming the existence of a set of orthogonal basis functions $\psi_{i}$ whose linear combinations can represent the signals, and since the signals themselves belong to the subspace spanned by these basis functions, it follows that each basis function can be represented as a linear combination of the signals. This is essentially a generalization of the Gram-Schmidt orthogonalization to function spaces, where a new basis function $\psi_{\text{new}}^{\text{gs}}(x)$ is constructed according to a new signal $u_{\text{new}}(x)$ as follows:
$\psi_{\text{new}}^{\text{gs}}(x)=u_{\text{new}}(x)-\sum_{j=1}^{|\Psi^{\text{gs}}|}\alpha^{\text{new}}_{j}\psi_{j}^{\text{gs}}(x),$ (16)
where the right-hand side is the residual between the new signal and its best reconstruction with the current basis functions available in the dictionary. Hence, the coefficients $\alpha^{\text{new}}_{j}$ are the projections of the new signal onto the previously found basis functions $\psi_{j}^{\text{gs}}(x)$, see Eq. (13). It is straightforward to show that if $u_{\text{new}}(x)$ does not belong to the dictionary subspace, the residual signal $\psi_{\text{new}}^{\text{gs}}(x)$ is orthogonal to the subspace. The complete procedure of constructing an appropriate dictionary with (weakly) orthogonal basis functions parameterized by neural networks is presented in Algorithm 2.
Algorithm 2 Dictionary Learning (Sample-wise)
1:Input:
2:$\mathcal{D}_{u}=\left\{\mathcal{D}_{u}^{(i)}\right\}_{i=1}^{N},\mathcal{D}_{u}^{(i)}=\left\{\left(\boldsymbol{x}^{(i,j)},\bar{u}^{(i,j)}\right)\right\}_{j=1}^{M^{(i)}}$triangleright $N$ realization of the input function $\bar{u}(\boldsymbol{x})$
3:$\boldsymbol{\Psi}^{\text{gs}}=\emptyset$triangleright A trivial dictionary
4:$\eta\in\mathbb{R}^{+}$triangleright Optimizer learning rate
[…]
6:Output: $\boldsymbol{\Psi}^{\text{gs}}=\left\{\psi_{l}^{\text{gs}}(\cdot;\boldsymbol{\theta}_{l})\right\}_{l=1}^{|\boldsymbol{\Psi}^{\text{gs}}|}$triangleright A parameterized dictionary by $|\boldsymbol{\Psi}^{\text{gs}}|$ neural networks
7:for $i\leftarrow 1$ to $N$ do
8:     $r^{(i)}(\boldsymbol{x};\boldsymbol{\alpha},\boldsymbol{\theta})=u^{(i)}(\boldsymbol{x})-{\boldsymbol{\Psi}^{\text{gs}}}^{T}(\boldsymbol{x};\boldsymbol{\theta})\ \boldsymbol{\alpha}$triangleright The residual function associated with the $i$-th signal
9:     $\boldsymbol{\alpha}^{(i)}=\underset{\boldsymbol{\alpha}}{\text{argmin}}\ \underset{\boldsymbol{x}\sim\mathcal{D}_{u}^{(i)}}{\mathbb{E}}\left\|r^{(i)}(\boldsymbol{x};\boldsymbol{\alpha},\boldsymbol{\theta})\right\|_{2}^{2}+\mathcal{R}(\mathcal{\boldsymbol{\alpha}})$triangleright The best estimator of basis coefficients
10:     $\text{Err}=\underset{\boldsymbol{x}\sim\mathcal{D}_{u}^{(i)}}{\mathbb{E}}\left\|r^{(i)}(\boldsymbol{x};\boldsymbol{\alpha}^{(i)},\boldsymbol{\theta})\right\|_{2}^{2}$triangleright The estimated residual norm of the best estimator
[…]
12:         continue
13:     end if
14:     $\boldsymbol{\theta}_{\text{new}}=\underset{\bar{\boldsymbol{\theta}}}{\text{argmin}}\ \underset{(\boldsymbol{x},\bar{u})\sim\mathcal{D}_{u}^{(i)}}{\mathbb{E}}\left\|\psi^{\text{gs}}_{\text{new}}(\boldsymbol{x};\bar{\boldsymbol{\theta}})-r^{(i)}(\boldsymbol{x};\boldsymbol{\alpha}^{(i)},\boldsymbol{\theta})\right\|_{2}^{2}$triangleright Train a new basis to capture the residual signal
15:     $\boldsymbol{\Psi}^{\text{gs}}\leftarrow\boldsymbol{\Psi}^{\text{gs}}\cup\{\psi^{\text{gs}}_{\text{new}}(\boldsymbol{x};\boldsymbol{\theta}_{\text{new}})\}$triangleright Update the dictionary
16:end for
Discussion: Computational Complexity and Efficiency – Although these two algorithms share similarities, they may have different computational costs and training stability. For example, time complexity of Algorithm 2 scales linearly with the number of realizations, which can be mitigated by the batch-wise training feature of Algorithm 1. Additionally, the basis functions identified by Algorithm 2 may be more sensitive to the order of the data stream since the first basis is derived from the first data it encounters. Moreover, when the data per realization is severely sparse, this algorithm may produce initial basis functions that are overfitted and not representative of the entire set of realizations. Consequently, under such scenarios, the initial basis functions may introduce highly oscillatory components, leading to noisy information during the basis identification for other realizations, which do not reflect the statistical properties of the entire dataset. The answer to all these questions requires a separate study and comparison between these two algorithms under various scenarios, which is beyond the main focus of this research. Hence, in this work, we primarily focus on the application of Algorithm 1 and only showcase the application of Algorithm 2 in the final numerical example.
Discussion: Normalization – From a theoretical perspective, the introduced algorithms aim to (weakly) enforce orthogonality of the identified basis functions. However, for several reasons, one might be interested in basis functions that are normalized as well, i.e., orthonormal basis functions. One reason could be attributed to potential identifiability issues associated with Eq. (11), where both the basis functions and coefficients are unknown. If one scales up the coefficients by a factor and simultaneously scales down the basis functions by the same factor, the same representation can be achieved, although the unknowns themselves have changed. The ADMM method mitigates this issue by solving for each unknown separately while fixing the others, thereby preventing simultaneous changes. However, controlling the scale or norm of either set of unknowns may further suppress such potential issues. Also, the relative scale of each basis function compared to others may impact the scale of its corresponding optimal coefficient relative to the coefficients of other basis functions. Since we may later use the basis coefficients in other downstream tasks (here, operator learning), it may be helpful to have learned bases with relatively the same scaling or norm. This way, the scale of their corresponding coefficients will have a particular meaning in terms of their relative importance in the prediction (coefficients with higher values indicate that their corresponding bases have more impact in the prediction). Several strategies are possible for the normalization of the learned basis functions: (1) one may use a normalization layer in the output layer of the neural network during the training to adjust the basis functions according to the batch statistics. In this study, we used an $L_{2}$-norm normalization layer of this type, where the normalized basis is $\psi_{\text{norm}}(\boldsymbol{x};\boldsymbol{\theta})=\psi(\boldsymbol{x};\boldsymbol{\theta})/\sqrt{\mathbb{E}_{x}[\psi^{2}(\boldsymbol{x};\boldsymbol{\theta})]}$. This type of normalization shares similarities with other batch-wise normalization layers commonly used in ML [1, 49]. (2) If the batch statistics of the given data do not accurately represent the actual statistics, one may normalize the learned basis after its training. To do this, a sufficiently large number of points can be artificially generated to empirically estimate the norm of the current un-normalized basis function, using a Markov Chain Monte Carlo (MCMC) estimate. This scaling factor can then be saved for later use before proceeding to learn the next basis function. In this work, we used the first approach in our numerical examples unless otherwise specified.
@@ 3 Numerical Examples
In this section, we showcase the resolution independence of RINO by solving four numerical examples, where the data is synthesized from PDE solutions under different setups, ranging from 1D operators to operators with different supports in their input and output functions. The setup for the data generation and operator learning task is specified for each problem separately in their respective sections. Examples in Sections 3.1.1, 3.1.2, 3.1.3 and 3.1.4 are chosen from the literature or designed to verify the resolution independence of RINO with respect to the input function data, which is the main focus of this study. Example in Section 3.2 is designed to demonstrate the application of RINO in a more challenging scenario where the data realizations are obtained from finite element simulations of varying resolutions. The compactness and consistency of the vectorial representation achieved through our dictionary learning algorithm are discussed and studied in comparison with random projections in Appendix  F.
In cases where the PDE solver for data generation works on a fixed discretization, we randomly select only a subset of input function sensors for training to showcase RINO’s ability to handle irregular data. The locations and numbers of these sensors vary from sample to sample. Assuming an input function realization is discretized at $M$ points, a corresponding “random" realization with random discretization is identified by randomly selecting $M_{\text{rand}}$ points without replacement, where $M_{\text{rand}}$ is independently drawn for each realization from the uniform distribution $M_{\text{rand}}\sim\text{Uniform}[M_{\text{min}},M_{\text{max}}]$, and $0<M_{\text{min}}<M_{\text{max}}\leq M$.
@@
3.1 Verification Examples
3.1.1 Example 1: Antiderivative
@@
where $u(x)$ is modeled as a Gaussian random process with zero mean and a covariance function defined by the radial basis function (RBF) with length scale $l=0.2$, i.e., $u(x)\sim\mathcal{GP}(0,\text{Cov}(x_{1},x_{2}))$; details are provided in Appendix G. We use the open-source dataset from [22], which includes 150 pairs of input-output functions for the training set and 1000 pairs for the test set. In the original dataset, the functions are discretized at 100 equally spaced sensor points. Following the procedure described earlier, we set $M_{\text{min}}=10$ and $M_{\text{max}}=60$ to generate random point clouds for the training and testing cases. In other words, each training data function has a randomly subsampled discretization of between 10 and 60 points.
@@
As described in the formulation section, the first step in RINO is to learn a dictionary to represent the input functions. This dictionary will be used to find embedding coordinates for the operator learning task, specifically for use in the branch network. Figure 4(a) shows the reconstruction errors in Algorithm 1 during the ADAM [19] iterations. Each red cross on the curve indicates the step at which a new basis function is added (line 8 in Algorithm 1). With a total of 9 learned basis functions, the data can be reconstructed with high accuracy, as shown by the distribution of errors in Figure 4(b) for both training and testing datasets. The learned basis functions are plotted in Figure 4(c).
@@
Figure 5: Antiderivative Example: (a) Loss function during optimization iterations for operator learning. (b) Distribution of output function prediction errors for the training and testing datasets after training.
@@
The ADAM iterations for operator learning using the input function embeddings obtained from the learned dictionary are shown in Figure 5(a). The close gap between the training and testing errors during the iterations indicates good generalization behavior, suggesting that the input function representation is informative. In Figure 5(b), the distribution of output function prediction errors after training is shown for both the training and testing sets.
@@
Figure 6: Antiderivative Example: RINO prediction for two queried input functions at inference and for the testing dataset: (a) Worst-case sample with the highest error in output function prediction, (b) Sample at the 25th percentile of the error distribution (i.e., 75% of cases have less error than this value). The first column shows the queried input function data as ‘o’ and the masked points (not provided to the dictionary) as ‘x’, along with the reconstruction from dictionary learning. The second column shows the projection coefficients of the queried input function data onto the learned dictionary of basis function from Figure 4(c). The last column shows the output function prediction along with the true solution.
@@
In Figure 6, the prediction of RINO at inference is shown for two queries from the test dataset. Figure 6(a) presents the results for the worst-case output function prediction. The first column shows the queried input function, which includes a point cloud with 35 random data points shown as ‘o’ along with the masked data points shown as ‘x’ that are not provided to the dictionary but are used to assess the accuracy of the reconstruction. The second column shows the projection coefficients of the input function data onto the learned dictionary, which has a total of 9 basis functions. The reconstruction using these coefficients is also plotted in the first figure. The last column shows the prediction of the output function along with the true model solution. Similarly, Figure 6(b) presents the same results for the test sample at the 25th percentile of the error distribution, meaning 75% of test cases have lower error than this sample. As the results show, the RINO method demonstrates good performance in providing predictions at inference for input function data recorded at an arbitrary (but sufficiently rich) number of sensors and locations.
3.1.2 Example 2: Nonlinear 1D Darcy’s Equation
@@
The worst-case RINO predictions are shown in Figure 7(a), demonstrating visually acceptable accuracy for both the output function prediction and the input function reconstruction. Similarly the test sample at the 25th percentile of the error distribution is shown in Figure 7(b).
Figure 7: 1D Nonlinear Darcy’s Equation: RINO Prediction for two queried input functions at inference and for the testing dataset: (a) Worst-case sample with the highest error in output function prediction, (b) Sample at the 25th percentile of the error distribution (i.e., 75% of cases have less error than this value). The first column shows the queried input function data as ‘o’ and the masked points (not provided to the dictionary) as ‘x’, along with the reconstruction from dictionary learning. The second column shows the projection coefficients of the queried input function data onto the learned dictionary of basis function from Figure 4(c). The last column shows the output function prediction along with the true solution.
3.1.3 Example 3: Nonlinear 2D Darcy’s Equation
@@
where the solution-dependent permeability is $\kappa(s(\boldsymbol{x}))=0.2+s^{2}(\boldsymbol{x})$ and the source term is a random field $u(\boldsymbol{x})\sim\mathcal{GP}$ with length scale $l=0.25$. Homogeneous Dirichlet boundary conditions $s=0$ are defined at the domain boundaries. The solution field is obtained via the finite element method. Input function random fields are discretized on a $20\times 20$ uniform grid, while the triangulation of the finite element domain does not necessarily coincide with these input sensor point locations. Following the procedure described earlier, we set $M_{\text{min}}=100$ and $M_{\text{max}}=280$ to generate random point clouds over a subset of the 400 grid points for the training and testing cases. The sizes of the training and test datasets are 800 and 200, respectively.
The RINO predictions corresponding to the worst-case test error for the output field is shown in Figure 8, demonstrating satisfactory error.
Figure 8: 2D Nonlinear Darcy’s Equation: RINO prediction for the worst-case output function $s(\boldsymbol{y})$ in the test dataset. The top left figure shows the queried input function data $u(\boldsymbol{x})$, with black circle symbols indicating the queried data. The middle top figure shows the input field reconstruction via the learned dictionary. The top right figure shows the error between the reconstructed input field and the ground truth. The bottom figures show the output field $s(\boldsymbol{y})$, the RINO prediction, and its corresponding errors.
3.1.4 Example 4: Burgers’ Equation
@@
where periodic boundary conditions are assumed, and the periodic initial condition is a Gaussian random process with parameters similar to [42], and $\nu=0.01$. A spectral solver is used to obtain the solution field on a $101\times 100$ grid in the space-time domain for each input function, which is discretized on 101 uniformly spaced points. Following the procedure described earlier, we set $M_{\text{min}}=40$ and $M_{\text{max}}=70$ to generate random point clouds for the training and testing cases. The sizes of the training and test datasets are 1500 and 500, respectively. Notice that the input function samples are randomly discretized; however, the predictions are on the full solution grid for the results in this section, similar to other examples.
@@
To showcase the flexibility of the RINO framework, we use predefined trunk basis vectors as the first 70 POD bases obtained from the solution output snapshots of the training dataset. It has been shown previously that using predefined optimal bases can facilitate the training of DeepONet and achieve faster convergence [23]. Two samples from the testing set are shown in Figure 9.
@@
Figure 9: 2D Burger’s Equation: RINO prediction for two queried initial condition fields at inference and for the testing dataset: (a) Sample with the highest error in output function prediction $s(\boldsymbol{y})$, and (b) Sample at the 25th percentile of the error distribution (i.e., 75% of cases have less error than this value). The first column shows the ground truth solution. The second column shows the RINO prediction with the queried input initial condition points shown along the x-axis. The last column shows the point-wise error. The output domain $\boldsymbol{y}=(x,t)$ has spatio-temporal coordinates, where the horizontal axis represents space and the vertical axis represents time.
3.1.5 Discussion
@@
After training each of the models listed above on the defined randomly sub-sampled data, we can perform inference during testing using any combination of input function sensors and at any desired points, thanks to the fully continuous nature of the RINO for processing both input and output functions. To examine the influence of discretization, we perform additional tests on regularly subsampled sensors for the input function at inference. A summary of our results is provided in Table 1. Rows in this table with a numeric value for the number of sensors (e.g., $M=100$) indicate performance during inference for input functions queried at structured and uniformly spaced sensor locations. Rows in the table denoted ‘random,’ correspond to test sets on the randomly sub-sampled discretizations described in Sections 3.1.1 – 3.1.4. The RINO used for predictions is trained using a data set of randomly sub-sampled points as described for each problem in Sections 3.1.1 – 3.1.4. Therefore, training error is reported only for the ‘random’ case.
@@
The error statistics shown in Table 1 are computed for the trained models described above. Each neural network model is randomly initialized before training, and the sensor locations for the input functions are also randomly generated for each training run.
@@
As the results suggest, the output function prediction is nearly insensitive to the discretization resolution of the queried input function for all PDE cases. However, in the 1D Nonlinear Darcy’s and Burgers’ equations, we observe a degradation in accuracy after some level of resolution coarsening (see the last row for each case). This occurs because, as anticipated, when the resolution falls below a certain threshold—particularly when it is significantly less than the problem’s characteristic length scale—the predictions become erroneous and thus resolution-dependent. In other words, these correspond to cases where a ReNO cannot be accurately identified due to under-sampling.
Number of Sensors ($M$) Test Error (mean $\pm$ std) Train Error
@@
Table 1: Resolution dependence study for each verification problem. RINO output function prediction errors are reported for queried input functions at various resolutions denoted by the Number of Sensors ($M$). Cases where a numerical value is prescribed correspond to fixed regular discretizations of the input. Cases denoted ‘random’ indicate a random subsampling of the input discretization as defined in Sections 3.1.1 – 3.1.4. Mean training/test errors and standard deviations are computed from five independent trials.
3.2 Demonstration Example: Nonlinear 2D Darcy’s Equation on Unstructured Meshes
In this problem, we revisit the 2D Darcy flow problem previously introduced in Section 3.1.3. The main difference is that, in this example, the data belongs to unstructured finite element meshes that differ from realization to realization. The goal here is not to showcase the method’s consistency across varying resolutions (from low to high or vice versa), which was studied in previous problems. Instead, we aim to demonstrate that the RINO can perform operator learning for problems of practical interest involving arbitrary finite element meshes.
To generate training data, we purposely generate unstructured triangular meshes with varying density per realization. The areas with higher mesh density are arbitrary and do not correspond, for example, to areas where the mesh might be refined based on an error estimator. We then interpolate the previously generated random input fields (see Section 3.1.3) onto the new unstructured meshes and obtain the solution field for each realization using a nonlinear finite element solver written in FEniCS [2].
Operator Learning with unknown trunk
Similar to the previous examples, using Algorithm 1, we learn 40 INR basis functions to effectively project an arbitrary input function point cloud onto the dictionary subspace with a satisfactory reconstruction error (see Appendix D). We then use this 40-dimensional latent representation for operator learning where the trunk and branch neural networks are unknown and are trained. The worst-case output function prediction and the case corresponding to the 25th percentile of the error distribution from the unseen test data are shown in  Figures 10 and 11, respectively. The results indicate a satisfactory training process and prediction capability. The full error distributions for the training and test data are provided in Appendix D.
Figure 10: 2D Nonlinear Darcy’s Equation (Unstructured Mesh): RINO prediction with the worst-case output function, $s(\boldsymbol{y})$, in the test dataset. The top left figure shows the queried input function data $u(\boldsymbol{x})$. The middle top figure shows the input field reconstruction via learned dictionary. The top right figure shows the error between the reconstructed input field and the ground truth. The bottom figures show the true output field $s(\boldsymbol{y})$, the RINO prediction, and its corresponding errors.
Figure 11: 2D Nonlinear Darcy’s Equation (Unstructured Mesh): RINO results for the test sample with an output function prediction error in the 25th percentile of the error distribution. The top left figure shows the queried input function data $u(\boldsymbol{x})$. The middle top figure shows the input field reconstruction via learned dictionary. The top right figure shows the error between the reconstructed input field and the ground truth. The bottom figures show the true output field $s(\boldsymbol{y})$, the RINO prediction, and its corresponding errors.
Operator Learning with predefined INR trunk
As discussed in Section 2.1.1, the proposed dictionary learning algorithms can be applied in a similar manner for learning appropriate basis functions for the output function data. Once the dictionary is learned (offline), it can be utilized as a predefined trunk. Under such a setup, we only need to train the branch network. In this problem, we utilize Algorithm 2 to learn a dictionary of 54 INR basis functions that can sufficiently approximate the output function data. The worst-case output function prediction and the case corresponding to the 25th percentile of the error distribution from the unseen test data are shown in  Figures 12 and 13, respectively. The results indicate a satisfactory training process and prediction capability. The full error distributions for the training and test data are provided in Appendix D.
Figure 12: 2D Nonlinear Darcy’s Equation (Unstructured Mesh) with Predefined INR Trunk Net: RINO prediction with the worst-case output function, $s(\boldsymbol{y})$, in the test dataset. The top left figure shows the queried input function data $u(\boldsymbol{x})$. The middle top figure shows the input field reconstruction via learned dictionary. The top right figure shows the error between the reconstructed input field and the ground truth. The bottom figures show the true output field $s(\boldsymbol{y})$, the RINO prediction, and its corresponding errors.
Figure 13: 2D Nonlinear Darcy’s Equation (Unstructured Mesh) with Predefined Trunk Net: RINO results for the test sample with an output function prediction error in the 25th percentile of the error distribution. The top left figure shows the queried input function data $u(\boldsymbol{x})$. The middle top figure shows the input field reconstruction via learned dictionary. The top right figure shows the error between the reconstructed input field and the ground truth. The bottom figures show the true output field $s(\boldsymbol{y})$, the RINO prediction, and its corresponding errors.